OpenStudy (anonymous):

can someone show me how to slove for n 6+3n=18

OpenStudy (cwrw238):

subtract 6 from both sides of the '=' then divide both sides by 3

OpenStudy (cwrw238):

6 + 3n = 18 -6 -6

OpenStudy (cwrw238):

6-6 = ? 18 -6 =?

OpenStudy (anonymous):

sorry for long reply 6-6 is 0 and 18-6 is 12

OpenStudy (anonymous):

so no solution?

OpenStudy (anonymous):

i got it sorry

OpenStudy (cwrw238):

no 6 - 6 + 3n = 18 - 6 3n = 12 now divideboth sides by 3 to get value of n

OpenStudy (anonymous):

4

OpenStudy (cwrw238):

right